# Multi-Modal Attention Networks with Uncertainty Quantification for Automated Concrete Bridge Deck Delamination Detection

### v1

@@ 4.4  Attention Head Configuration
We evaluate how the number of attention heads affects performance. Table 5 reveals an interesting pattern: validation accuracy and F1-score remain invariant across configurations (79.09% and 0.294 respectively), while AUC improves substantially from 0.469 (hh=1) to 0.611 (hh=4), with marginal additional gains to 0.613 (hh=8). This divergence indicates that multi-head attention primarily enhances confidence calibration rather than altering decision boundaries. The training dynamics differ markedly: single-head attention exhibits slow convergence (6.8% accuracy at epoch 1) with negative train-validation gap indicating underfitting, while eight-head attention achieves 60.8% at epoch 1 with healthy positive gap. These findings suggest four to eight heads suffice for multi-modal NDE fusion, with larger counts providing minimal benefits.
@@
4.5  Per-Bridge Performance Analysis
The most important findings emerge from per-bridge evaluation, which reveals how performance depends on dataset characteristics. Tables 610 present comprehensive comparisons across all five bridges at six learning rates for traditional machine learning methods, Simple CNN, and the proposed attention-based approach.
On the Park River Median dataset (Table 6), where class distribution is relatively balanced (approximately 56% Class 1, 33% Class 2, 11% Class 3), the proposed attention-based method achieves strong performance: 96.6% accuracy, 99.8% AUC, and 96.3% F1-score at learning rate $5\times 10^{-4}$. This represents improvements of approximately 18–20 percentage points in accuracy and 24 percentage points in F1-score over the Simple CNN baseline, which achieves 77.2% accuracy and 72.0% F1-score at the same learning rate. Traditional machine learning methods perform less strongly, with SVM and Random Forest achieving around 71–72% accuracy.
The temporal attention module processes GPR A-scans through multi-head self-attention following 1D convolutional feature extraction, enabling the network to selectively focus on diagnostic temporal patterns within the 512-sample radar signals. The spatial attention module for IRT images combines channel attention and spatial attention mechanisms to identify thermally anomalous regions. The cross-modal fusion layer leverages multi-head attention to learn complementary relationships between GPR and IRT modalities, dynamically weighting the more reliable modality for each sample. This adaptive fusion explains the substantial AUC improvements (99.8% vs 83.2% for Simple CNN), as the model learns optimal probabilistic rankings by integrating evidence from both sensors.
@@ 4.5  Per-Bridge Performance Analysis
Similarly strong results emerge on the Park River North Bound dataset (Table 7), where despite a class distribution of approximately 78% Class 1, 22% Class 2, and 0.4% Class 3, the attention mechanism achieves 95.5% accuracy and 99.5% AUC compared to Simple CNN’s 93.2% accuracy and 92.0% AUC. The architectural advantage manifests particularly in the model’s ability to maintain relatively high minority class detection rates. The attention mechanisms’ selective weighting allows the network to develop specialized feature representations for the rare Class 3 samples by learning which temporal patterns in GPR signals and which spatial patterns in thermal images most reliably indicate this critical defect type.
@@
The Forest River Median North Bound dataset (Table 8) provides additional evidence of attention’s utility under moderate imbalance. Despite a 61:28:10 class distribution, the proposed method achieves 73.5% accuracy and 92.5% AUC, outperforming Simple CNN by approximately 10 percentage points in accuracy (63.5% vs 73.5%). When training data includes sufficient minority class samples (approximately 10% representation), the multi-head attention layers learn discriminative weightings, with different heads potentially specializing in different defect characteristics. This head specialization enables the network to capture the multi-faceted nature of delamination signatures that manifest differently across GPR and IRT modalities depending on defect depth, extent, and moisture content.
@@
While the attention method achieves lower raw accuracy, it demonstrates substantially better AUC (92.5% vs. 75.0%), suggesting superior probabilistic ranking of predictions even when final classification decisions are less accurate. This divergence between AUC and accuracy indicates that the attention mechanism’s learned feature representations maintain meaningful internal structure—the temporal attention, spatial attention, and cross-modal fusion components successfully learn discriminative patterns that separate classes in probability space—but the final classification layer struggles to set appropriate decision thresholds under emerging class imbalance. The F1-score reveals a concerning pattern: the proposed method achieves only 74.5% compared to Simple CNN’s 75.4%. This near-parity despite the large AUC advantage hints at early signs of the challenges that become more pronounced in severely imbalanced datasets.
@@
The attention model’s 24.8 percentage point gap between accuracy (81.5%) and F1-score (56.7%) contrasts with Simple CNN’s minimal 3.8 percentage point gap (91.8% accuracy vs. 88.0% F1). This pattern suggests that while the attention mechanism achieves reasonable overall accuracy, it does so primarily by correctly predicting the dominant class while struggling to maintain balanced performance across minority classes. We hypothesize that the temporal attention heads, designed to identify diagnostic patterns across GPR signals, instead learn to recognize variations within normal concrete responses because 89.5% of training gradients flow from Class 1 samples. Similarly, the spatial attention mechanism may learn to focus on patterns characteristic of sound bridge decks rather than thermal signatures of delamination. With only approximately 2,000 Class 3 samples among over 100,000 total samples, the attention weights receive negligible gradient updates encouraging recognition of deep delamination patterns.
Traditional machine learning methods actually outperform the attention-based approach on this dataset, with SVM achieving 89.6% accuracy and 84.7% F1-score. This finding suggests that when class imbalance becomes severe, approaches with simpler decision boundaries—which lack the capacity to develop elaborate within-class feature hierarchies—may be more robust than complex deep architectures with attention mechanisms.
@@
4.6  Discussion
The per-bridge analysis reveals that attention-based fusion substantially outperforms baselines when class imbalance ratios remain below approximately 8:1, achieving 10–25 percentage point improvements in accuracy and AUC on four of five bridges. Under extreme imbalance exceeding 9:1 with minority classes below 2%, the architectural properties enabling these advantages—learned feature weighting, multi-head specialization—appear to become liabilities. The increased model capacity that benefits learning on balanced data enables overfitting to majority class patterns when minority classes are severely underrepresented.
@@ 5  Conclusion
This research addressed automated bridge deck delamination detection through a multi-modal attention network that fuses Ground Penetrating Radar and Infrared Thermography data with integrated uncertainty quantification. The deteriorating state of civil infrastructure—with over 231,000 U.S. bridges exhibiting structural deficiencies—demands inspection techniques that overcome the limitations of single-modal approaches, where GPR struggles with shallow defects and IRT lacks depth penetration.
@@
The proposed architecture introduces three synergistic innovations: temporal attention for GPR signal processing that selectively focuses on diagnostic reflection patterns, spatial attention for IRT imagery that identifies thermally anomalous regions, and cross-modal attention fusion with learnable modality embeddings that discovers complementary defect patterns across sensing modalities. Comprehensive uncertainty quantification decomposes predictive uncertainty into epistemic and aleatoric components, enabling selective prediction strategies that achieve 93.2% accuracy by rejecting uncertain cases for human review.
@@
Our systematic experimental investigation across five SDNET2021 bridge datasets yielded nuanced insights about when attention-based fusion delivers advantages. On four datasets with balanced to moderately imbalanced class distributions, the architecture achieves substantial improvements: 10–25 percentage point gains in accuracy and AUC compared to single-modal approaches and simple fusion baselines. Park River Median exemplified these strengths with 96.6% accuracy and 99.8% AUC. Ablation studies confirmed that cross-modal attention provides critical gains (+2.8%) beyond within-modality attention alone, while multi-head mechanisms achieve 30% improvement in AUC through enhanced confidence calibration. Uncertainty quantification reduces Expected Calibration Error by 72%, enabling reliable confidence estimates essential for safety-critical deployment.
@@
However, the investigation also characterized important boundary conditions. On Park River South Bound, exhibiting extreme class imbalance (89.5:8.7:1.8), the attention mechanism achieved only 56.7% F1-score compared to Simple CNN’s 88.0%. This vulnerability appears to stem from insufficient gradient signal from minority classes, causing attention heads to optimize for majority class variations rather than developing specialized minority class detectors. Notably, this represents one of five evaluated datasets—the architecture performs favorably across typical inspection scenarios.
@@
These findings provide actionable deployment guidance. Attention-based fusion represents the preferred approach for bridge inspection datasets with class imbalance ratios below approximately 8:1, where its sophisticated cross-modal learning capabilities translate to meaningful performance advantages. For the minority of cases exhibiting extreme imbalance (minority classes below 2%), practitioners should consider simpler baselines or implement specialized techniques such as focal loss or class reweighting. The architecture’s compact parameterization (14.8M parameters) and practical inference time (28.4ms) enable real-time bridge deck scanning during field inspections.
@@
Several directions merit future investigation. Cross-dataset generalization and transfer learning capabilities remain unexplored—whether models trained on one bridge can successfully inspect another has significant practical implications. Systematic investigation of whether focal loss, SMOTE oversampling, or two-stage training can restore attention mechanism advantages on severely imbalanced datasets would extend the method’s applicability. Alternative attention formulations, including hierarchical or sparse attention mechanisms, might exhibit different sensitivity profiles worthy of exploration.
@@
In conclusion, this work demonstrates that multi-modal attention networks with uncertainty quantification represent a genuine advance for automated bridge infrastructure inspection, consistently outperforming traditional approaches across diverse real-world conditions while providing well-characterized failure modes that enable informed deployment decisions. By delivering both strong performance on typical scenarios and honest characterization of boundary conditions, this research contributes actionable knowledge for deploying automated inspection systems that combine machine efficiency with appropriate human oversight for safety-critical infrastructure management.

### v3

@@ 4.3  Multi-Head Attention Configuration Analysis
We investigate how attention head count affects performance characteristics. Table 2 reveals an interesting divergence: validation accuracy and F1-score remain constant across configurations (79.09% and 0.294 respectively), while AUC improves substantially from 0.469 (hh=1) to 0.611 (hh=4), with marginal additional gains to 0.613 (hh=8). This divergence indicates that multi-head attention primarily enhances confidence calibration rather than modifying decision boundaries. Training dynamics differ markedly: single-head attention exhibits slow convergence (6.8% accuracy at epoch 1) with negative train-validation gap indicating underfitting, while eight-head attention achieves 60.8% at epoch 1 with healthy positive gap. These findings suggest four to eight heads suffice for multi-modal NDE fusion, with larger counts providing diminishing returns.
@@
4.4  Individual Bridge Performance Characterization
The most significant findings emerge from per-bridge evaluation, revealing how performance depends on dataset characteristics. Tables 37 present comprehensive comparisons across all five bridges at six learning rates for traditional machine learning methods, Simple CNN, and the proposed attention-based approach.
On the Park River Median dataset (Table 3), where class distribution is relatively balanced (approximately 56% Class 1, 33% Class 2, 11% Class 3), the proposed attention-based method demonstrates strong performance: 96.6% accuracy, 99.8% AUC, and 96.3% F1-score at learning rate $5\times 10^{-4}$. This represents improvements of approximately 18–20 percentage points in accuracy and 24 percentage points in F1-score over the Simple CNN baseline, which achieves 77.2% accuracy and 72.0% F1-score at the same learning rate. Traditional machine learning methods perform less favorably, with SVM and Random Forest achieving approximately 71–72% accuracy.
The temporal attention module processes GPR A-scans through multi-head self-attention following 1D convolutional feature extraction, enabling selective focus on diagnostic temporal patterns within the 512-sample radar signals. The spatial attention module for IRT images integrates channel attention and spatial attention mechanisms to identify thermally anomalous regions. The cross-modal fusion layer employs multi-head attention to learn complementary relationships between GPR and IRT modalities, dynamically weighting the more reliable modality for each sample. This adaptive fusion explains the substantial AUC improvements (99.8% vs 83.2% for Simple CNN), as the model learns optimal probabilistic rankings by integrating evidence from both sensors.
@@ 4.4  Individual Bridge Performance Characterization
Similarly strong results emerge on the Park River North Bound dataset (Table 4), where despite a class distribution of approximately 78% Class 1, 22% Class 2, and 0.4% Class 3, the attention mechanism achieves 95.5% accuracy and 99.5% AUC compared to Simple CNN’s 93.2% accuracy and 92.0% AUC. The architectural advantage manifests particularly in the model’s capacity to maintain relatively high minority class detection rates. The attention mechanisms’ selective weighting enables the network to develop specialized feature representations for rare Class 3 samples by learning which temporal patterns in GPR signals and which spatial patterns in thermal images most reliably indicate this critical defect type.
@@
The Forest River Median North Bound dataset (Table 5) provides additional evidence of attention’s utility under moderate imbalance. Despite a 61:28:10 class distribution, the proposed method achieves 73.5% accuracy and 92.5% AUC, outperforming Simple CNN by approximately 10 percentage points in accuracy (63.5% vs 73.5%). When training data includes sufficient minority class representation (approximately 10%), the multi-head attention layers learn discriminative weightings, with different heads potentially specializing in different defect characteristics. This head specialization enables the network to capture the multi-faceted nature of delamination signatures that manifest differently across GPR and IRT modalities depending on defect depth, extent, and moisture content.
@@
While the attention method achieves lower raw accuracy, it demonstrates substantially superior AUC (92.5% vs. 75.0%), suggesting better probabilistic ranking of predictions even when final classification decisions are less accurate. This divergence between AUC and accuracy indicates that the attention mechanism’s learned feature representations maintain meaningful internal structure—the temporal attention, spatial attention, and cross-modal fusion components successfully learn discriminative patterns that separate classes in probability space—but the final classification layer struggles to establish appropriate decision thresholds under emerging class imbalance. The F1-score reveals a concerning pattern: the proposed method achieves only 74.5% compared to Simple CNN’s 75.4%. This near-parity despite the large AUC advantage hints at early signs of the challenges that become more pronounced in severely imbalanced datasets.
@@
The attention model’s 24.8 percentage point gap between accuracy (81.5%) and F1-score (56.7%) contrasts with Simple CNN’s minimal 3.8 percentage point gap (91.8% accuracy vs. 88.0% F1). This pattern suggests that while the attention mechanism achieves reasonable overall accuracy, it does so primarily by correctly predicting the dominant class while struggling to maintain balanced performance across minority classes. We hypothesize that the temporal attention heads, designed to identify diagnostic patterns across GPR signals, instead learn to recognize variations within normal concrete responses because 89.5% of training gradients originate from Class 1 samples. Similarly, the spatial attention mechanism may learn to focus on patterns characteristic of sound bridge decks rather than thermal signatures of delamination. With only approximately 2,000 Class 3 samples among over 100,000 total samples, the attention weights receive negligible gradient updates encouraging recognition of deep delamination patterns.
Traditional machine learning methods actually outperform the attention-based approach on this dataset, with SVM achieving 89.6% accuracy and 84.7% F1-score. This finding suggests that when class imbalance becomes severe, approaches with simpler decision boundaries—which lack the capacity to develop elaborate within-class feature hierarchies—may prove more robust than complex deep architectures with attention mechanisms.
@@
4.5  Synthesis and Practical Implications
The per-bridge analysis reveals that attention-based fusion substantially outperforms baselines when class imbalance ratios remain below approximately 8:1, achieving 10–25 percentage point improvements in accuracy and AUC on four of five bridges. Under extreme imbalance exceeding 9:1 with minority classes below 2%, the architectural properties enabling these advantages—learned feature weighting, multi-head specialization—appear to become vulnerabilities. The increased model capacity that benefits learning on balanced data enables overfitting to majority class patterns when minority classes are severely underrepresented.
@@ 5  Conclusion
We have presented a comprehensive investigation of multi-modal attention networks integrating Ground Penetrating Radar and Infrared Thermography data for automated bridge deck delamination detection with uncertainty quantification. The investigation was motivated by the critical infrastructure challenge facing civil systems worldwide—with over 231,000 U.S. bridges exhibiting structural deficiencies—requiring inspection methodologies that transcend the limitations of single-modality approaches, where GPR lacks sensitivity to shallow defects and IRT cannot penetrate to deeper structural layers.
@@
Our architectural contribution integrates three complementary attention mechanisms: temporal self-attention for GPR signal processing that selectively emphasizes diagnostic electromagnetic reflection patterns, spatial attention for IRT imagery that localizes thermally anomalous regions, and cross-modal attention fusion employing learnable modality embeddings that discovers complementary defect patterns across sensing technologies. Integrated Bayesian uncertainty quantification decomposes predictive uncertainty into epistemic and aleatoric components, facilitating selective prediction strategies that achieve 93.2% accuracy by routing uncertain cases to human expert review.
@@
Systematic experimental investigation across five SDNET2021 bridge datasets yielded nuanced insights regarding the conditions under which attention-based fusion delivers performance advantages. On four datasets exhibiting balanced to moderately imbalanced class distributions, the architecture achieves substantial gains: 10–25 percentage point improvements in accuracy and AUC relative to single-modality approaches and concatenation-based fusion baselines. The Park River Median dataset exemplified these strengths with 96.6% accuracy and 99.8% AUC. Comprehensive ablation studies confirmed that cross-modal attention provides critical gains (+2.8%) beyond within-modality attention operating independently, while multi-head configurations achieve 30% improvement in AUC through enhanced confidence calibration. Uncertainty quantification reduces Expected Calibration Error by 72%, enabling reliable confidence estimates essential for safety-critical deployment scenarios.
@@
However, our investigation also characterized important boundary conditions where attention-based approaches face challenges. On Park River South Bound, exhibiting extreme class imbalance (89.5:8.7:1.8), the attention mechanism achieved only 56.7% F1-score compared to Simple CNN’s 88.0%. This vulnerability appears to originate from insufficient gradient signal from minority classes, causing attention heads to optimize for majority class pattern variations rather than developing specialized minority class detectors. Importantly, this represents one of five evaluated datasets—the architecture performs favorably across typical inspection scenarios encountered in practice.
@@
These findings provide actionable deployment guidance for practitioners. Attention-based fusion represents the recommended approach for bridge inspection datasets exhibiting class imbalance ratios below approximately 8:1, where its sophisticated cross-modal learning capabilities translate to meaningful performance advantages. For the minority of cases exhibiting extreme imbalance (minority classes below 2%), practitioners should consider simpler architectural baselines or implement specialized techniques including focal loss or class reweighting strategies. The architecture’s efficient parameterization (14.8M parameters) and practical inference latency (28.4ms) enable real-time bridge deck scanning during field inspection operations.
@@
Several directions warrant future investigation. Cross-dataset generalization and transfer learning capabilities remain unexplored—whether models trained on one bridge structure can successfully inspect different bridges has significant practical implications for deployment scalability. Systematic investigation of whether focal loss, SMOTE oversampling, or two-stage training protocols can restore attention mechanism advantages on severely imbalanced datasets would extend the method’s applicability range. Alternative attention formulations, including hierarchical or sparse attention mechanisms, might exhibit different sensitivity profiles worthy of systematic exploration.
@@
In summary, this work demonstrates that multi-modal attention networks with integrated uncertainty quantification represent a genuine advancement for automated bridge infrastructure inspection, consistently outperforming traditional approaches across diverse real-world conditions while providing well-characterized failure modes that enable informed deployment decisions. By delivering both strong performance on typical scenarios and honest characterization of boundary conditions, this investigation contributes actionable knowledge for deploying automated inspection systems that combine machine efficiency with appropriate human oversight for safety-critical infrastructure management.